You have discovered a field where archery was done several years ago. The find includes a full lane with the line from which arrows are shot, and a target with an arrow still stuck in it. You measure the distance from the line to the target to be 27.4 meters. The arrow has landed at an angle of 4 degrees below the horizontal. To help confirm if this is a legitimate find, calculate the veolocty at which the arrow was launched. Assume the initital velocity toward the target is entirely horizontal, and the angle of the arrow in the target is the direction of final velocity
Answer #1

Vertical displacement is given by : h = d tan theta = 27.4*tan 4 degree =  1.916 m

Using the second equation of motion, h = 0.5gt^2

t = sqrt(2h/g) = sqrt(2*27.4/9.8)

= 2.365 s

initial velocity v = distance/time = 27.4m/2.365s

= 11.6 m/s answer